what should be the width of each slit to obtain 10 maxima of the double slit pattern within the central maximum of the single slit pattern of wavelength 500 nm and separation between slits is 1mm

Grade:

2 Answers

Arun
5 years ago

the slit width-wavelength relation is given as

 

aθ = λ

where 

a is the slit width

θ is the angular separation  and

λ is the wavelength

 

so,

θ = λ/a

 

now, in this case we have

width of  each slit = width of ten maxima

so,

10(λ/d) = 2(λ/a)

or

a = d/5

thus, the slit width will be

a = 1/5 = 0.2 mm

Pragya Yadav
3 years ago
We want a theta=lemda,theta=lemda/a                      10 lemda/d =2 lemda/a or, a=d/5=0.2 mm. Notice that the wavelength of light and distance of the screen do not enter in the calculation of a.